Future values.Number of Periods Future Value Present Value $282.00 $17,476.00 $37,595.00 $27,545.00 Interest Rate 4% 7.5% 11% 16% 26

Fill in the future values for the following​ table, using one of the three methods​ below:

a.  Use the future value​ formula,

FV equals PV times left parenthesis 1 plus r right parenthesis Superscript nFV=PV×(1+r)n.

b.  Use the TVM keys from a calculator.

c.  Use the TVM function in a spreadsheet.
Answer #1

Using the formula

1.
=282*(1+4%)^6
=356.8199632

2.
=17476*(1+7.5%)^30
=153001.5969

3.
=37595*(1+11%)^26
=566927.518

4.
=27545*(1+16%)^14
=220016.183